Liquid water is to be compressed by a pump whose isentropic efficiency is 81 percent from 0.2 MPa to 5 MPa at a rate of 0.15Answer: 14.8 kW, Explain?
Answer #1

Given P 0.2 MPa = 200 kPa P2 = 5 MPa = 5000 kPa Flow rate, Q = 0.15 m3/min = (0.15/60) m3/s = 0.0025 m3/s Isentropic efficien

For an isentropic process ds = 0. Therefore dhis = vdP Expressing this in finite difference form Ahis = v(P2 - P.) Substituti

Wact 12 0.81 14.81 kW Power input to the compressor = 14.8 kW

MatLab code to compute the power

1 - clear all close all clc coo) un WN 5 - Q = 0.15/60; P2 = 5*10^3; P1 = 0.2*10^3; eta = 0.81; flow rate, m^3/s $ Final pres

MatLab output

Command Window Power input to the compressor: 14.815 kW